Hi i am finding it hard to answer this question for my homework the question is:
Find the solution of the equation 2x - 9 = 25
Thank you hope you can help

Answer by dolly(163)

2x-9=25
Adding 9 to both the sides of the equation,
2x-9+9=25+9
2x=34
2x/2=34/2
x=17